OpenStudy (anonymous):

|3x+5| = 11 Solve for x

hero (hero):

3x + 5 = 11 3x+ 5 = -11

hero (hero):

jahtoday, there's actually a solution for this one

OpenStudy (anonymous):

3x+5=11 or 3x+5=-11 Because it is absolute value, it can be positive or negative. 3x=6 or 3x=-16 x=2 or -16/3

hero (hero):

Careful...absolute value can only be positive. The negative should be understood as moving to the left from a certain number

OpenStudy (anonymous):

If you plug the answers in, you will get... |3(2)+5| = |11| = 11 and |3(-16/3)+5| = |-16+5| = |-11| = 11

hero (hero):

I didn't say your answer was wrong, I just said be careful how you explain it.